# InterACT: Inter-dependency Aware Action Chunking with Hierarchical Attention Transformers for Bimanual Manipulation

### v2

@@ Abstract
We present InterACT: Inter-dependency aware Action Chunking with Hierarchical Attention Transformers, a novel imitation learning framework for bimanual manipulation that integrates hierarchical attention to capture inter-dependencies between dual-arm joint states and visual inputs. InterACT consists of a Hierarchical Attention Encoder and a Multi-arm Decoder, both designed to enhance information aggregation and coordination. The encoder processes multi-modal inputs through segment-wise and cross-segment attention mechanisms, while the decoder leverages synchronization blocks to refine individual action predictions, providing the counterpart’s prediction as context. Our experiments on a variety of simulated and real-world bimanual manipulation tasks demonstrate that InterACT significantly outperforms existing methods. Detailed ablation studies validate the contributions of key components of our work, including the impact of CLS tokens, cross-segment encoders, and synchronization blocks.
Keywords: Imitation Learning, Bimanual Manipulation
@@ 1 Introduction
Bimanual manipulation tasks, such as unscrewing a bottle cap or connecting two electrical cables, presents significant challenges due to the need for high precision coordination. Traditional approaches often rely on high-end robots and precise sensors, which can be expensive and require meticulous calibration [1, 2, 3]. However, recent advances in learning-based approaches offer the potential to perform such complex tasks using low-cost hardware.
The ALOHA and the Action Chunking with Transformers (ACT) framework has shown that low-cost systems can achieve high-precision tasks that were traditionally only possible with expensive setups. The ACT addresses the compounding error problem in imitation learning by predicting sequences of actions rather than single steps, thereby reducing the task’s effective horizon and mitigating errors over time [4].
Despite these recent advances, it remains a challenge for bimanual robotics to ensure robust and accurate coordination between two arms in a dynamic environment. In this work, we propose InterACT: a new policy for bimanual manipulation that emphasizes inter-dependencies between two arms by utilizing hierarchical attention mechanisms. In our designs, multimodal inputs are encoded through segment-wise and cross-segment encoders which handle the complex relationships between different segments in a manner similar to how long documents are processed in NLP [5]. This combines the proprioceptive data of the robot arm joints and the visual features of the camera in a coherent latent space that allows for coordinated detail-oriented and smooth action execution.
Our method goes beyond existing approaches by focusing on the hierarchical nature of bimanual tasks, ensuring each arm operates independently while coordinating seamlessly with its counterpart and the environment. Building on the latest advances in attention mechanisms, this work presents a compelling solution and features improved performance in bimanual manipulation tasks.
Contributions:
1.
Hierarchical Attention Mechanism for Bimanual Manipulation: We adapt a novel hierarchical attention mechanism that effectively captures inter-dependencies between dual-arm joint states and visual inputs. This mechanism processes multi-modal inputs through segment-wise encoding followed by cross-segment encoding, which captures inter-dependencies and allows for precise and coordinated bimanual manipulation.
2.
Comprehensive Evaluation: We conduct extensive experiments on multiple simulated and real-world implemented tasks. The results highlight the effectiveness of our approach compared to previous methods.
3.
Ablation Studies: We perform several ablation studies to assess contributions from different components of our framework, namely the classification (CLS) tokens, the cross-segment encoder, and the synchronization block. These studies provide insight into the importance of each element to realize bimanual manipulation that is robust and efficient.
@@ 2.1 Bimanual Manipulation
Bimanual manipulation involves two robotic arms performing tasks that require dexterity and synchronization. Inspired by the natural ability of humans to perform such tasks, researchers have been keen on modeling these skills in robots. Prevailing methodologies, including classical control methods, reinforcement learning, and imitation learning, have significantly advanced the field.
Early research primarily relied on classical control methods, focusing on predefined trajectories and high-fidelity models to achieve coordinated movements [1, 2, 3]. However, these methods often required extensive calibration and were less adaptable to dynamic environments.
The introduction of reinforcement learning (RL) made bimanual manipulation more adaptable and robust. RL-based approaches have proven effective in handling complex tasks, outperforming classical methods, and generalizing across different scenarios [6, 7, 8, 9]. These methods leverage RL’s ability to learn from interactions with the environment, improving performance in varied and unpredictable conditions.
Imitation learning has emerged as a prominent method for teaching robots bimanual tasks through human demonstrations. This approach enables robots to mimic complex human actions, facilitating the execution of intricate tasks. Research has demonstrated the effectiveness of imitation learning in training robots for coordinated dual-arm manipulation using hierarchical skill learning and force-based techniques [10, 11, 12, 13, 14].
Recent advancements have focused on integrating machine learning models to enhance imitation learning. Frameworks like ALOHA have shown that low-cost systems can perform high-precision tasks using imitation learning techniques traditionally reserved for expensive setups [4, 15]. Similarly, the Action Chunking with Transformers (ACT) algorithm addresses the compounding error problem in imitation learning by predicting sequences of actions, improving accuracy and efficiency [4].
Despite these advancements, bimanual manipulation remains challenging, especially in dynamic and unstructured environments. Recent research has explored integrating additional multi-modal data, such as language or sensory feedback, to enhance the robustness and efficiency of bimanual manipulation systems [16, 17]. These efforts hold promise for developing more adaptable and efficient robotic systems capable of performing a wide range of complex manipulation tasks.
@@ 2.2 Hierarchical Attention Mechanisms
In robotic manipulation, the concept of hierarchy has been explored to manage the complexity of long-horizon tasks by breaking down problems into manageable sub-tasks [25]. As proven in long document classification, leveraging segment-wise and cross-segment attention mechanisms, hierarchical attention models can capture dependencies within and across segments [5]. This makes hierarchical attention transformers particularly appealing for use in bimanual robotic manipulation tasks, where precise coordination between arms is crucial.
In this research, we explore a different utility of hierarchical attention in which inter/intra-segment attention supports a more robust extraction of inter-dependencies between the actions of the two arms in handling a single task, leading to more coordinated actions in complex bimanual manipulation.
@@ 3 InterACT: Inter-dependency Aware Action Chunking with Hierarchical Attention Transformer
InterACT builds upon the Action Chunking with Transformers (ACT) framework [4], enhancing it with hierarchical attention mechanisms to capture inter-dependencies between dual-arm joint states and visual inputs. This section details the components and workflow of the InterACT framework, focusing on the encoder and decoder structures within an imitation learning framework.
The ACT framework leverages transformer architecture to predict future steps in bimanual manipulation tasks, effectively handling sequences of actions by capturing temporal dependencies. However, it does not explicitly model inter-dependencies between dual-arm joint states and visual inputs, which can limit its performance in complex manipulation tasks.
InterACT extends the ACT framework by introducing hierarchical attention transformers, consisting of two main components: Hierarchical Attention Encoder and Multi-arm Decoder.
@@ 3.1 Hierarchical Attention Encoder
Segments are defined as individual groups of data inputs that are processed independently before being integrated. In this context, segments include the joint states of each arm and visual features at a specific timestep. Hierarchical Attention Encoder processes input segments and captures both intra-segment and inter-segment dependencies through a hierarchical attention mechanism, which consists of segment-wise encoder and cross-segment encoder layers. A detailed pipeline for the Hierarchical Attention Encoder is illustrated in Figure 1 and Algorithm 1.
Input: Each joint in each arm is embedded into a single token through a linear layer. Visual features are extracted from RGB images using ResNet18 backbones, which convert and flatten the images along the spatial dimension to form a sequence of feature tokens. The joint embeddings from both arms and the visual embeddings of image frames from multiple cameras are concatenated to form a combined sequence. Multiple classification (CLS) tokens are prepended to each segment, allowing the model to summarize segment information during attention processing. Positional embeddings are then added to the tokens to retain sequence information, ensuring the model can understand the order of the tokens within each segment. The input for each decoder includes relevant segments and CLS tokens to facilitate cross-attention mechanisms.
Segment-wise Encoder: Each segment passes through multiple layers of multihead self-attention to aggregate information within each segment. The self-attention mechanism allows each token within the segment to attend to every other token, including the CLS tokens, thereby capturing intra-segment dependencies. This aggregation of information enables the model to effectively leverage the relationships within different parts of the segment [5], such as the connections between joints within the same arm and the relationships within the features from visual inputs. Importantly, the segment-wise encoder uses shared weights across all segments in each layer, ensuring uniform transformation and processing.
Cross-segment Encoder: The Cross-segment Encoder handles CLS tokens from each segment to capture inter-segment relationships. This component allows the model to discern and combine information across different segments [5] (e.g., joint states of both arms and visual features). The cross-segment encoder employs shared weights across all segments to maintain consistent modeling of these relationships.
Output: The original visual feature (input to the encoder, excluding CLS tokens) replaces the processed visual features to retain detailed spatial-temporal information.
@@ 3.2 Multi-arm Decoder
Similar to multi-task learning in NLP [26, 27], the Multi-arm Decoder comprises two parallel paths of decoder blocks, each dedicated to processing the encoded states and target tokens to generate the predicted actions for one of the arms. This section details the components and workflow of the Multi-arm Decoder, highlighting how it employs enriched embeddings from the encoder to coordinate actions between both arms effectively. A detailed pipeline for the Multi-arm Decoder is illustrated in Figure 2 and Algorithm 2.
Input: Encoded states from the Hierarchical Attention Encoder are used as the context for decoding. Target tokens, initialized with positional embeddings, serve as the starting point for generating actions. The input for each decoder includes relevant segments and CLS tokens to facilitate cross-attention mechanisms.
Arm Specific Decoders: The Arm Specific Decoders are responsible for generating intermediate decoder outputs for each arm. The Arm1 Specific Decoder takes in Arm1 segments, Arm2 CLS tokens, and visual information, while the Arm2 Specific Decoder takes in Arm2 segments, Arm1 CLS tokens, and visual information. Each layer processes these inputs through a cross-attention mechanism with the target tokens, incorporating contextual information from both joint states and visual features. This design enables each arm to use contextual information from the other arm ensuring synchronized actions.
Synchronization Block: To enhance coordination between both arms, before the final layer of each decoders, the intermediate outputs from the two decoders are concatenated and processed through a synchronization block, which uses self-attention to integrate the shared information. This step ensures that both arms leverage the combined context from the other arm and visual inputs before the final decoder block. The use of attention mechanisms for sharing information across multiple decoders has also been explored in multi-task learning, demonstrating its effectiveness in improving performance and coherence within tasks [28]. The output of the synchronization block is then input to the arm specific decoders for the final outputs.
@@ 4 Experiments and Results
For the real-robot setup, we modified the ALOHA 2 [29] setup by adjusting the height of the top camera to improve the visibility of the tabletop environment. This adjustment ensures that the camera captures a more comprehensive view of the workspace, which is essential for accurately tracking the bimanual manipulation tasks. Our robot setup is illustrated in Appendix C.
To evaluate our model, we conducted experiments on three simulation tasks and six real-world tasks: Transfer Cube and Peg Insertion along with Slide Ziploc and Thread Velcro are tasks adapted from ACT [4]. We introduce five new tasks: one simulation task and four real-world tasks. The simulation task is Slot Insertion, where both arms need to grab each side of a long peg together and place it in a slot on the table. The new real-world tasks include Insert Plug, Click Pen, Sweep, and Unscrew Cap. Detailed task definitions are provided in Appendix A.
We focus our comparisons exclusively on ACT as ACT already outperforms BC-ConvMLP [30, 31], BeT [32], and VINN [33] by a large margin in bimanual manipulation tasks [4].
@@ 4.1 Results
The results of our experiments are summarized in Tables 1 and 2. Our InterACT model shows superior performance compared to ACT on all simulated and real-world tasks. In the simulated tasks, InterACT outperformed ACT significantly, particularly in the Transfer Cube and Peg Insertion tasks where coordination and precision are crucial. The success rates for the ”Transfer” stages in the Transfer Cube task, as well as the ”Insert” stages in the Peg Insertion task, were notably higher with InterACT, demonstrating the effectiveness of our method in tasks that require coordination between the two arms. Moreover, in newly introduced tasks such as Slot Insertion and Insert Plug, InterACT also demonstrated higher performance. The Slot Insertion task, which requires precise coordination between two arms to carry the peg and adjust for alignment, showed a 100% success rate with InterACT, compared to 88% with ACT. Similarly, in the Insert Plug task, InterACT achieved better results in the coordination subtask ”Insert”. This highlights the robustness of our model in handling tasks that require precise coordination between the two arms.
@@
4.2 Ablation Studies
In this section, we perform ablation studies to evaluate the contributions of different components of the InterACT framework. Specifically, we focus on the impact of CLS tokens, the cross-segment encoder, and the synchronization block in the decoder. Detailed results of these ablation studies are provided in Appendix B.
@@ 4.2 Ablation Studies
| InterACT (Ours) | 98 | 88 | 84 | 88 | 78 | 44 | 100 | 100 |
Impact of Synchronization Block: The synchronization block enhances coordination between the two arms by sharing contextual information during decoding. This is crucial for synchronized and efficient bimanual manipulation, especially for predicting a sequence of actions. The results show that the removal of the synchronization block leads to a significant drop in performance across all tasks, particularly in the Transfer Cube and Peg Insertion tasks. This demonstrates the necessity of the synchronization block for achieving coordinated and synchronized actions.
The ablation studies clearly illustrate that all components of the proposed InterACT framework—CLS tokens, cross-segment encoder, and synchronization block—play a critical role in achieving high success rates in coordination tasks. The highest performance for the final coordination task is achieved when all components are utilized, underscoring the importance of the holistic integration of these elements in the InterACT framework.

### v3

@@ Abstract
Bimanual manipulation presents unique challenges compared to unimanual tasks due to the complexity of coordinating two robotic arms. In this paper, we introduce InterACT: Inter-dependency aware Action Chunking with Hierarchical Attention Transformers, a novel imitation learning framework designed specifically for bimanual manipulation. InterACT leverages hierarchical attention mechanisms to effectively capture inter-dependencies between dual-arm joint states and visual inputs. The framework comprises a Hierarchical Attention Encoder, which processes multi-modal inputs through segment-wise and cross-segment attention mechanisms, and a Multi-arm Decoder that generates each arm’s action predictions in parallel, while sharing information between the arms through synchronization blocks by providing the other arm’s intermediate output as context. Our experiments, conducted on various simulated and real-world bimanual manipulation tasks, demonstrate that InterACT outperforms existing methods. Detailed ablation studies further validate the significance of key components, including the impact of CLS tokens, cross-segment encoders, and synchronization blocks on task performance. We provide supplementary materials and videos on our project page111https://soltanilara.github.io/interact/.
Keywords: Bimanual Manipulation, Imitation Learning
@@ 1 Introduction
Bimanual manipulation tasks, such as unscrewing a bottle cap or connecting two electrical cables, presents significant challenges due to the need for high precision and complex coordination between the two arms. Traditional approaches often rely on high-end robots and precise sensors, which can be expensive and require meticulous calibration [1, 2, 3]. However, recent advances in learning-based approaches offer the potential to perform such complex tasks using low-cost hardware. The ALOHA and the Action Chunking with Transformers (ACT) framework has shown that low-cost systems can achieve high-precision tasks that were traditionally only possible with expensive setups. The ACT addresses compounding error problem in imitation learning by predicting sequences of actions rather than single steps, thereby reducing the task’s effective horizon and mitigating errors over time [4].
Despite recent advances, achieving the complex coordination between both arms required for consistent and successful execution of even relatively simple tasks remains a challenge in bimanual robotics. In this work, we propose InterACT: a new policy designed for bimanual manipulation that emphasizes inter-dependencies between two arms by utilizing hierarchical attention mechanisms. In our designs, multimodal inputs are encoded through segment-wise and cross-segment encoders which handle the complex relationships between different segments in a manner similar to how long documents are processed in NLP [5]. This combines the proprioceptive data of the robot arm joints and the visual features of the camera in a coherent latent space that allows for coordinated detail-oriented and smooth action execution.
Our method advances beyond existing approaches by emphasizing the hierarchical structure of bimanual tasks: while independent arm movements are managed at the lower level, coordination occurs at a higher level, allowing each arm’s actions to be informed by the other’s movements as well as the environmental sensory feedback.. Leveraging the latest advances in attention mechanisms, this work presents a novel solution for bimanual manipulation tasks.
@@ 2.1 Bimanual Manipulation
Bimanual manipulation involves two robotic arms performing tasks that require dexterity and coordination. Inspired by the natural ability of humans to perform such tasks, researchers have been keen on modeling these skills in robots. Prevailing methodologies, including classical control methods, reinforcement learning, and imitation learning, have significantly advanced the field.
Early research primarily relied on classical control methods, focusing on predefined trajectories and high-fidelity models to achieve coordinated movements [1, 2, 3]. However, these methods often required extensive calibration and were less adaptable to dynamic environments. The introduction of reinforcement learning (RL) made bimanual manipulation more adaptable and robust. RL-based approaches have proven effective in handling complex tasks, outperforming classical methods, and generalizing across different scenarios [6, 7, 8, 6, 9]. These methods leverage RL’s ability to learn from interactions with the environment, improving performance in varied and unpredictable conditions [10, 11, 12]. Imitation learning has emerged as a prominent method for teaching robots tasks through human demonstrations. This approach enables robots to mimic complex human actions, facilitating the execution of intricate tasks [13, 14]. Research has demonstrated the effectiveness of imitation learning in training robots for coordinated dual-arm manipulation leveraging waypoints, hierarchical skill learning and force-based techniques [15, 16, 17, 18, 19, 20]. Frameworks like ALOHA have shown that low-cost systems can perform high-precision tasks using imitation learning techniques traditionally reserved for expensive setups [4, 21]. Similarly, the Action Chunking with Transformers (ACT) algorithm addresses the compounding error problem in imitation learning by predicting sequences of actions, improving accuracy and efficiency [4].
Despite these advancements, bimanual manipulation remains challenging. Recent research has explored stabilizing one arm [22], simplifying actions [23], integrating additional multi-modal data, such as language or sensory feedback [24, 25, 26], to enhance the robustness and efficiency of bimanual manipulation systems. These efforts hold promise for developing more adaptable robotic systems capable of performing more complex manipulation tasks.
@@ 2.2 Hierarchical Attention Mechanisms
In robotic manipulation, the concept of hierarchy has been explored to manage the complexity of long-horizon tasks by breaking down problems into manageable sub-tasks [34], but not at the policy’s attention layer level. As proven in long document classification, leveraging segment-wise and cross-segment attention mechanisms, hierarchical attention models can capture dependencies within and across sentences [5]. This makes hierarchical attention transformers particularly appealing for use in bimanual robotic manipulation tasks, where capturing dependencies between arms is crucial.
In this work, we tailor the transformer architecture featuring hierarchical attention to the bimanual robotics tasks and explore its utility in extracting the complex inter-dependencies between the actions of the two arms. We hypothesize that this can lead to more coordinated actions and hence, more robust performance.
@@ 3 InterACT: Inter-dependency Aware Action Chunking with Hierarchical Attention Transformer
The ACT model [4] leverages transformer architecture to predict future steps in bimanual manipulation tasks, effectively handling sequences of actions by capturing temporal dependencies. However, it does not explicitly model inter-dependencies between dual-arm joint states and visual inputs, which can limit its performance in complex manipulation tasks.
InterACT builds upon the ACT model, enhancing it with hierarchical attention mechanisms to capture inter-dependencies between dual-arm joint states and visual inputs. This section provides an overview of the InterACT model and its key components: the Hierarchical Attention Encoder, which processes multi-modal inputs to capture both intra- (corresponding to one arm or sensory input) and inter-segment (across arms or sensory inputs) dependencies, and the Multi-arm Decoder, which generates synchronized action predictions for both arms.
@@ 3.1 Hierarchical Attention Encoder
Segments are defined as independent groups of data inputs that are processed individually before integration. In this context, segments include the joint states of each arm and visual features at a specific timestep. The Hierarchical Attention Encoder processes input segments and captures both intra-segment and inter-segment dependencies through a hierarchical attention mechanism that consists of two primary components: the Segment-wise encoder and the Cross-segment encoder. A detailed pipeline for the Hierarchical Attention Encoder is illustrated in Figure 1 and Algorithm 1.
Each raw joint position is embedded into a single token through a linear layer. Visual features are extracted from RGB images using ResNet18 backbones, which convert and flatten the images along the spatial dimension to form a sequence of visual feature tokens. The joint sequences from both arms and the visual feature sequence are then concatenated to form a combined input sequence. Classification tokens (CLS tokens) are prepended to each segment, allowing the model to capture and summarize segment information during attention [33, 35]. Positional embeddings are applied to the sequence to ensure the model can understand the order of the tokens within each segment sequence.
Segment-Wise encoder is responsible for capturing intra-segment dependencies by processing each segment individually. Each segment represents a distinct modality, and the encoder’s task is to aggregate the information within each segment. The self-attention mechanism allows each token within the segment to attend to every other token, including the CLS tokens, thereby allowing the CLS token to capture intra-segment dependencies. This aggregation of information enables the model to effectively leverage the relationships within different parts of the segment [5], such as the dependencies between joints or the spatial relationships within visual inputs.
Cross-segment encoder is responsible for capturing inter-segment dependencies by handling the CLS tokens generated from each segment by the segment-wise encoder. The CLS tokens serve as a condensed representation of each segment, summarizing the key information within the segment. By focusing on these CLS tokens, the cross-segment encoder reduces the computational overhead while effectively integrating information across different modalities.
Multiple segment-wise and cross-segment encoders are stacked, allowing the model to progressively refine its understanding of both intra- and inter-segment relationships. We follow an interleaved stacking approach [5], where segment-wise and cross-segment encoders alternate. This deep stacking enables the model to capture the complex dependencies across segments.
@@ 3.2 Multi-arm Decoder
Similar to multi-task learning in NLP [36, 37], the Multi-arm Decoder comprises two parallel paths of decoder blocks, each dedicated to processing the encoded states and target tokens to generate the predicted actions for one of the arms. This separation makes sense in bimanual manipulation tasks because, while the arms must coordinate, they often perform independent actions that require individual attention and control. By using separate decoders, each arm can process its specific actions while still sharing critical information through synchronization mechanisms.
This section details the components and workflow of the Multi-arm Decoder, highlighting how it employs enriched input sequence and CLS tokens from the encoder to coordinate actions between both arms effectively. A detailed pipeline for the Multi-arm Decoder is illustrated in Figure 1 and Algorithm 2.
Encoded states from the Hierarchical Attention Encoder are used as the context for decoding. Target tokens, initialized with positional embeddings, serve as the starting point for generating sequence of actions. The input for each decoder includes relevant segments and CLS tokens to facilitate cross-attention mechanisms.
Arm Specific Decoders are responsible for generating intermediate decoder outputs for each arm. The Arm1 Decoder takes in Arm1 segments, Arm2 CLS tokens, and visual feature segments, while the Arm2 Decoder takes in Arm2 segments, Arm1 CLS tokens, and visual feature segments. Each layer processes these inputs through a cross-attention mechanism with the target tokens, incorporating contextual information from both joint states and visual features. This design enables each arm to use contextual information from the other arm ensuring synchronized actions.
Synchronization Block enhances coordination between both arms. Before generating the final output from each decoder, the intermediate outputs from both decoders are concatenated and processed through a synchronization block, which uses self-attention to integrate shared information. This step ensures that both arms leverage the combined context from the other arm and visual inputs before passing through the reset of the decoder layers. The use of attention mechanisms for sharing information across multiple decoders has also been explored in multi-task learning, demonstrating its effectiveness in improving performance and coherence within tasks [38].
@@ 4 Experiments and Results
For the real-robot setup, we modified the ALOHA 2 [39] setup by adjusting the height of the top camera to improve the visibility of the tabletop environment. This adjustment ensures that the camera captures a more comprehensive view of the workspace, which is essential for accurately tracking the bimanual manipulation tasks. Our robot setup is illustrated in Appendix B.
To evaluate our model, we conducted experiments on three simulation tasks and six real-world tasks: Transfer Cube and Peg Insertion along with Slide Ziploc and Thread Velcro are tasks adapted from ACT [4]. We also introduce five new tasks: one simulation task and four real-world tasks. The simulation task is Slot Insertion, where both arms need to grab each side of a long peg together and place it in a slot on the table. The new real-world tasks include Insert Plug, Click Pen, Sweep, and Unscrew Cap. We collected 50 demonstrations for each task. For the simulation tasks, the data used to train the model were all from human demonstrations and we did not evaluate performance of the models trained on scripted data. Detailed task definitions are provided in Appendix A, and the hyperparameters of both ACT and InterACT are provided in Appendix C.
To Evaluate the performance of InterACT, we chose to compare our model against ACT [4] and the Diffusion policy [40]. However, with only 50 demonstrations, the Diffusion policy failed to execute any of the nine tasks. We did not perform any additional tuning on the Diffusion policy, as its poor performance suggested it was not well-suited for tasks with limited demonstration data. Therefore, we focus our comparisons exclusively on ACT, as ACT already significantly outperforms several other policies, including BC-ConvMLP [41, 14], BeT [42], and VINN [43], by a large margin in bimanual manipulation tasks [4].
@@ 4.1 Results
The results of our experiments are summarized in Tables 1. Our InterACT model shows superior performance compared to ACT on all simulated and real-world tasks. In the simulated tasks, InterACT outperformed ACT significantly, particularly in the Transfer Cube and Peg Insertion tasks where coordination and precision are crucial. The success rates for the ”Transfer” stages in the Transfer Cube task, as well as the ”Insert” stages in the Peg Insertion task, were notably higher with InterACT, demonstrating the effectiveness of our method in tasks that require coordination between the two arms. Moreover, in newly introduced tasks such as Slot Insertion and Insert Plug, InterACT also demonstrated higher performance. The Slot Insertion task, which requires precise coordination between two arms to carry the peg and adjust for alignment, showed a 100% success rate with InterACT, compared to 88% with ACT. Similarly, in the Insert Plug task, InterACT achieved better results in the coordination subtask ”Insert”. This highlights the robustness of our model in handling tasks that require precise coordination between the two arms.
@@
| InterACT (all components) | 98 | 88 | 84 | 88 | 78 | 44 | 100 | 100 |
4.2 Ablation Studies
In this section, we perform ablation studies to evaluate the contributions of different components of the InterACT framework. Specifically, we focus on the impact of CLS tokens, the cross-segment encoder, and the synchronization block in the decoder.
To gain deeper insights into the model’s behavior, we studied how attention weights to CLS tokens at the decoder change over the timesteps. The results illustrated in Figure 2 showed that during the phase of interaction between the two arms, significant spikes in attention weights were observed. These spikes occurred at key moments where coordinated actions between the arms were necessary. This indicates that the model heavily relies on the CLS tokens to process information when coordinating actions between the arms. This observation highlights the CLS tokens’ importance in facilitating precise bimanual manipulation.
Impact of Synchronization Block: The synchronization block enhances coordination between the two arms by sharing contextual information at the decoder.. This is crucial for synchronized and efficient bimanual manipulation. The results show that the removal of the synchronization block leads to a significant drop in performance across all tasks, particularly in the Transfer Cube and Peg Insertion tasks. This demonstrates the necessity of the synchronization block for achieving coordinated and actions.
Our ablation studies clearly illustrate that all components of the proposed InterACT framework—CLS tokens, cross-segment encoder, and synchronization block—play a critical role in achieving high success rates in coordination tasks. The best performance for the coordination subtask is achieved when all components are utilized, underscoring the importance of the holistic integration of these elements in the InterACT framework.